The weights of students in a certain school are normally distributed with a mean weight of 66 kg. 10% have a weight greater than 70kg. What percentage of students weighs between 52kg and 66kg?  


How did we get our answer for this problem above? 



1
Expert's answer
2022-04-13T15:05:31-0400

"P(X>70)=0.1,\\
P(X<70)=1-P(X>70)=\\=1-0.1=0.9."

From z-table the 90th percentile

is z = 1.282.

So,

"z=\cfrac{x-\mu} {\sigma}, \\
1.282=\cfrac{70-66}{\sigma},\\
\sigma=\cfrac{4} {1.282} =3.12,\\
z_1=\cfrac{52-66} {3.12}=-4.49, \\
z_2=\cfrac{66-66}{3.12}=0,\\"

"P(52<X<66) =P(-4.49<Z<0)=\\
=P(Z<0)-P(Z<-4.49)=\\=0.5-0=0.5=50\%\\\text{ (from z-table). }"


We found the standard deviation and then the searched probability.